# Jacobian Scopes: token-level causal attributions in LLMs

### v1

@@ A.5 Path-Integrated Semantic Scope
$\mathrm{Influence}_{t}:=\left\lVert\mathrm{IG}(\bm{X})_{t}\right\rVert_{2},$ (10)
@@
Fig.˜14 compares the resulting IG attribution with the original Semantic Scope on the example discussed in Section˜3.1. Relative to Semantic Scope, the IG-based method assigns disproportionately large influence to early-context tokens, thereby obscuring the semantically decisive input token “Columbia.”
@@
To diagnose the origin of this behavior, we visualize the integrand $\nabla z(\bm{\tilde{X}}(\alpha))$ at several values of $\alpha$ in Fig.˜15. Specifically, for each $\alpha$ we plot the token-wise $\ell_{2}$ norm of the gradient, $\|\nabla_{\bm{x}_{t}}z(\bm{\tilde{X}}(\alpha))\|_{2}$, consistent with the influence definition in Eq.˜10. In practice, the integral in Eq.˜9 is approximated by a discrete sum over 100 uniformly spaced values of $\alpha$; the figure shows a subset of these values.
@@
Crucially, the gradient magnitudes in this low-$\alpha$ regime are also large, causing these out-of-distribution effects to dominate the path integral in Eq.˜9. At intermediate values of $\alpha$, the attribution is often highest for function words such as “and,” which are not semantically relevant to the prediction. Taken together, these observations indicate that the IG path spends substantial measure in regions of representation space that are both out-of-distribution and attributionally pathological.
@@
A.6 Geometric interpretation of influence under $\varepsilon$ -norm perturbations
@@
A.7 Second-order KL expansion and Fisher pullback
@@
For a categorical distribution parameterized by logits $\bm{z}\in\mathbb{R}^{|\mathcal{V}|}$ with $\bm{p}=\mathrm{softmax}(\bm{z})$, the Fisher information with respect to $\bm{z}$ takes the closed form Bishop and Nasrabadi (2006)
@@
$\bm{F}_{u}$ $=\bm{W}^{\intercal}\bigl(\mathrm{Diag}(\bm{p})-\bm{p}\bm{p}^{\intercal}\bigr)\bm{W}.$
A.8 Expected KL under isotropic unit-norm perturbations
Consider the local quadratic expansion
$\mathrm{KL}\!\left(\bm{p}(\cdot\mid\bm{x})\,\middle\|\,\bm{p}(\cdot\mid\bm{x}+\bm{\delta})\right)$ $=\frac{1}{2}\,\bm{\delta}^{\intercal}\bm{F}_{t}\,\bm{\delta}+o(\|\bm{\delta}\|^{2})$
and isotropic unit-norm perturbations $\bm{\delta}=\varepsilon\bm{u}$, where $\varepsilon>0$ is small and $\bm{u}\in\mathbb{R}^{d}$ is uniform on the unit sphere ($\|\bm{u}\|_{2}=1$). Taking expectation and using $\mathbb{E}_{\bm{u}}[\bm{u}\bm{u}^{\intercal}]=\tfrac{1}{d}\bm{I}$ yields
@@ A.8 Expected KL under isotropic unit-norm perturbations
Thus, up to the universal factor $\varepsilon^{2}/(2d)$, the trace $\mathrm{tr}(\bm{F}_{t})$ governs the expected KL divergence induced by an isotropic unit-norm perturbation of the input.

### v2

@@ A.7 Path-Integrated Semantic Scope
$\mathrm{Influence}_{t}^{\mathrm{IG}}:=\left\lVert\mathrm{IG}(\bm{X})_{t}\right\rVert_{2},$ (12)
@@
Fig.˜13 compares the resulting IG attribution with the original Semantic Scope on the example discussed in Section˜3.1. Relative to Semantic Scope, the IG-based method assigns disproportionately large influence to early-context tokens, thereby obscuring the semantically decisive input token “Columbia.”
@@
To diagnose the origin of this behavior, we visualize the integrand $\nabla z(\bm{\tilde{X}}(\alpha))$ at several values of $\alpha$ in Fig.˜14. Specifically, for each $\alpha$ we plot the token-wise $\ell_{2}$ norm of the gradient, $\|\nabla_{\bm{x}_{t}}z(\bm{\tilde{X}}(\alpha))\|_{2}$, consistent with the influence definition in Eq.˜12. In practice, the integral in Eq.˜11 is approximated by a discrete sum over 100 uniformly spaced values of $\alpha$; the figure shows a subset of these values.
@@
Crucially, the gradient magnitudes in this low-$\alpha$ regime are also large, causing these out-of-distribution effects to dominate the path integral in Eq.˜11. At intermediate values of $\alpha$, the attribution is often highest for function words such as “and,” which are not semantically relevant to the prediction. Taken together, these observations indicate that the IG path spends substantial measure in regions of representation space that are both out-of-distribution and attributionally pathological.
@@
A.8 Geometric interpretation of influence under $\varepsilon$ -norm perturbations
@@
A.9 Second-order KL expansion and Fisher pullback
@@
For a categorical distribution $\bm{p}=\mathrm{softmax}(\bm{z})$, parameterized by logits $\bm{z}\in\mathbb{R}^{|\mathcal{V}|}$, its Fisher information with respect to $\bm{z}$ takes the closed form Bishop and Nasrabadi (2006)
@@
$\bm{F}$ $=\bm{W}^{\intercal}\bigl(\mathrm{Diag}(\bm{p})-\bm{p}\bm{p}^{\intercal}\bigr)\bm{W}.$
A.10 Fisher Scope as a low-rank approximation of total information
The Fisher Scope influence score in Eq.˜5 is based on the principal Fisher direction $\bm{u}_{1}$, the single direction in output space most sensitive to changes in the predictive distribution. Here we show that this can be understood as the leading term of a more general quantity, which aggregates distributional sensitivity across all directions.
Total information as expected KL divergence. To assess how an input token $\bm{x}_{t}$ influences the model’s prediction, we pull back Arvanitidis et al. (2022) the Fisher information matrix through the Jacobian $\bm{J}_{t}$ to the embedding space of token $t$: $\bm{F}_{t}:=\bm{J}_{t}^{\intercal}\bm{F}\bm{J}_{t}$. Consider an isotropic unit-norm perturbation $\bm{\delta}=\varepsilon\bm{u}$, where $\varepsilon>0$ is small and $\bm{u}\in\mathbb{R}^{d}$ is uniform on the unit sphere. Taking expectation and using $\mathbb{E}_{\bm{u}}[\bm{u}\bm{u}^{\intercal}]=\tfrac{1}{d}\bm{I}$ yields
@@ A.10 Fisher Scope as a low-rank approximation of total information
Thus, up to the universal factor $\varepsilon^{2}/(2d)$, the trace $\mathrm{tr}(\bm{F}_{t})$ governs the expected KL divergence induced by an isotropic perturbation of the input token $\bm{x}_{t}$. We therefore refer to $\mathrm{tr}(\bm{F}_{t})$ as the total information at position $t$.
Fisher Scope as rank-1 approximation. Since $\bm{F}$ admits the eigen-decomposition $\bm{F}=\bm{U}\bm{\Lambda}\bm{U}^{\intercal}$, the total information can be written exactly as
$\mathrm{tr}(\bm{F}_{t})=\sum_{i=1}^{d_{\mathrm{model}}}\lambda_{i}\|\bm{u}_{i}^{\intercal}\bm{J}_{t}\|_{2}^{2}.$ (14)
In practice, $\bm{F}$ admits a good low-rank approximation $\bm{F}\approx\bm{U}_{r}\bm{\Lambda}_{r}\bm{U}_{r}^{\intercal}$, where $\bm{U}_{r}\in\mathbb{R}^{d_{\mathrm{model}}\times r}$ contains the top $r$ eigenvectors and $\bm{\Lambda}_{r}$ the corresponding eigenvalues. This yields the approximation
$\mathrm{tr}(\bm{F}_{t})\approx\sum_{i=1}^{r}\lambda_{i}\|\bm{u}_{i}^{\intercal}\bm{J}_{t}\|_{2}^{2},$ (15)
where each term $\|\bm{u}_{i}^{\intercal}\bm{J}_{t}\|_{2}$ requires only a single backward pass, analogous to Eq.˜2. The full total information thus requires $r$ backward passes, compared to $d_{\mathrm{model}}$ passes needed to form $\bm{J}_{t}$ explicitly.
The Fisher Scope influence score in Eq.˜5 corresponds to the $r=1$ case of Eq.˜15, retaining only the leading term $\lambda_{1}\|\bm{u}_{1}^{\intercal}\bm{J}_{t}\|_{2}^{2}$ and dropping the scalar $\lambda_{1}$ since it is constant across token positions $t$.
Low-rankness of $\bm{F}$. Figure˜15 provides three complementary pieces of evidence that the $r=1$ approximation is justified in practice.
[…]
The top panel shows the eigenvalue spectrum of $\bm{F}$ across 10 LAMBADA prompts on a log-log scale: a consistent power-law decay spanning 6 orders of magnitude, indicating that the leading eigenvalue dominates all subsequent ones.
The middle panel confirms this quantitatively via the cumulative explained variance $\sum_{i=1}^{r}\lambda_{i}/\mathrm{tr}(\bm{F})$, averaged over 1000 LAMBADA passages: at $r=1$ the mean explained fraction is already $\approx 37\%$, rising to $\approx 65\%$ by $r=10$ and approaching 1 only beyond $r\approx 1000$.
The bottom panel provides direct empirical justification, extending the LLaMA-3.2 1B results in Table˜3: rank-1 Fisher Scope ($r=1$) already outperforms Input $\times$ Gradient and Semantic Scope on LAMBADA, and matches Temperature Scope. Increasing the rank to $r\in\{4,16,64,256\}$ yields only marginal further gains in AOPC, while multiplying the computational cost by $r$.
Together, these results justify both the privileged status of $\bm{u}_{1}$ and the choice of $r=1$ in Fisher Scope.